OpenStudy (anonymous):

Estimate the limit by substituting smaller and smaller values of h. Round the answer to one decimal place!! :) lim (e^4+h) - (e^4) h->0 -------------- h so trying to sub 0.1,0.01,0.001, and 0.000001 (for more reference hehe) do these look right? h=0.1... 57.42137 h=0.01...54.87205 h=0.001...54.625458 h=0.000001...54.598176 so would it be approaching 54 point something? not sure if that's right so far though. what do you think @zepdrix ?? :)

zepdrix (zepdrix):

h is in in the exponent position? e^(4+h) ?

zepdrix (zepdrix):

ya they look good!

OpenStudy (anonymous):

yes:)

zepdrix (zepdrix):

54 point something, 1 decimal again?

OpenStudy (anonymous):

okay yay!! :) but not sure what the decimal point would be :/ haha yeah, one decimal point:)

zepdrix (zepdrix):

The expression represents the limit definition of the derivative of the function e^x evaluated at x=4. The derivative of e^x is e^x. Evaluated at x=4 gives us e^4. So if you're having trouble figuring out exactly what value it's approaching, punch that into your calculator real quick. I know I know, it's kinda cheating.. but it's good to see the connection with these things.

OpenStudy (anonymous):

e^4=54.5981..... so does that give us the answer then? hahaha :P would it be 54.6 ?

zepdrix (zepdrix):

It's approaching 54.5981... So if you need to round to 1 decimal value, yah 54.6 is probably a good idea :O

OpenStudy (anonymous):

hehe okie yay:) thanks!!....again!! haha :P